You can afford monthly deposits of $170 into an account that pays 2.7% compounded monthly. How long until you have $9,100 to buy a boat?

Type the number of months: ___

(Please square, label, or circle the answer)
Answer #1
  • ANSWER IS 50 MONTHS

IN THIS QUESTION THE RATE OF INTEREST IS 2.7% COMPOUNDED MONTHLY

DEPOSIT AMOUNT IS $170 FOR A PARTICULAR PERIOD OF TIME

GIVEN $9100 AS THE FINAL AMOUNT

FORMULAE FOR CALCULATING COMPOUND INTEREST IS

             A=P(1+r/n)tn

A IS THE FINAL AMOUNT ,WHICH IS GIVEN IN THE QUESTION= 9100

P IS THE PRINCIPLE AMOUNT = 170

r IS THE RATE OF INTEREST = 2.7%        (2.7/100=0.027)

n   IS THE NUMBER OF TIMES WE DEPOSIT $170, HERE IT IS 12 ,BECAUSE MONTHLY DEPOSIT............

IF IT WAS YEARLY DEPOSIT, WE CAN TAKE n AS 1

T    IS THE TIME PERIOD OF DEPOSIT, WE HAVE TO FIND OUT THAT T

SO

9100=170(1+0.027/12)12*t

BY SOLVING THIS WE CAN GET 50 MONTHS AS THE ANSWER

THANK YOU